### v2

# Ab initio spin Hamiltonians and magnetism of Ce and Yb triangular-lattice compounds

Abstract
We calculate the crystal-field splitting, ground-state Kramers doublet and intersite exchange interactions within the ground-state doublet manifold using an ab initio Hubbard-I based approach for a representative set of Ce and Yb triangular-lattice compounds. These include the putative quantum spin liquids (QSL) RbCeO2 and YbZn2GaO5 and the antiferromagnets KCeO2 and KCeS2. The calculated nearest-neighbor (NN) couplings are antiferromagnetic and exhibit noticeable anisotropy. The next-nearest-neighbor (NNN) couplings are ferromagnetic in the Ce systems and dominated by classical dipole–dipole interactions in the Yb case. Solving the resulting effective spin-1/2 models by exact diagonalization up to $N=36$ sites, we predict ordered magnetic ground states for all systems, including the two QSL candidates. We explore the phase space of an anisotropic NN + isotropic NNN triangular-lattice model finding that a significant antiferromagnetic NNN coupling is required to stabilize QSL phases, while the NN exchange anisotropy is detrimental to them. Our findings highlight a possibly important role of deviations from the perfect triangular model – like atomic disorder – in real triangular-lattice materials.
Introduction. Triangular lattice spin-$1/2$ antiferromagnets (AFM), with their inherent geometric frustration and quantum fluctuations, stand as prime candidates for harboring exotic states of quantum matter, such as quantum spin liquids (QSL) and topologically ordered phases. While the nearest-neighbor Heisenberg model on the triangular lattice has been shown to stabilize a $120^{\circ}$ magnetically ordered Néel ground-state [1, 2, 3, 4], the inclusion of additional interactions has revealed a variety of both magnetically ordered and quantum disordered phases. Notably, enhanced second nearest-neighbor interactions can induce a quantum paramagnetic state governed by the physics of the gapless $U(1)$ Dirac spin liquid (DSL) [5, 6, 7], where distinct proposals have suggested a direct emergence of the DSL [8, 9, 10], a valence bond solid [11], a gapped $\mathbf{Z}_{2}$ spin liquids [12, 13, 14, 15] and chiral spin liquids [16, 17, 18], which are understood to be descendants of the parent DSL order [6, 7, 19, 20]. While the precise nature of the paramagnetic regime is still under debate, it has become evident that the physics of the DSL and its descendants serves as an organizing principle of the phase diagrams of triangular lattice antiferromagnets [11].
Geometrically perfect quasi-2D triangular lattices of magnetic ions are realized in numerous layered rare-earth oxides and chalcogenides. In particular, Kramers 3+ ions of RR=Ce, Yb, Nd etc. in systems like delafossites $ARX_{2}$ (where $A$ is an alkali metal, XX=O, S, Se) [21, 22, 23, 24, 25, 26, 27], heptatantalates RRTa7O19 [28], YbMgGaO4 [29, 30, 31, 32], and YbZn2GaO5 [33] have recently attracted a lot of interest. The crystal-field (CF) splitting in these systems is typically much larger than rather weak superexchange between the well localized 4ff shells of rare-earth ions. Consequently, admixture of the excited CF levels by superexchange can be neglected with the lowest Kramers doublet considered as an effective spin-1/2. However, strong spin-orbit (SO) entanglement in the 4ff ground state multiplet in conjunction with the weaker CF effect results in anisotropic charge and magnetic density of the Kramers states, which leads, in turn, to anisotropic superexchange and $g$-factors. Hence, the intersite coupling anisotropy that has been shown to strongly affect QSL formation is naturally present in those compounds. A gapless QSL ground state has been suggested on the basis of intensive experimental investigation for NaYbO2 [21], several Yb delafossite selenides and sulfides [34, 35] as well as for YbZn2GaO5 [33].
Low-energy Hamiltonians for those rare-earth triangular materials remain to date quite uncertain. In particular, due to lack of single crystals for majority of delafossites [33], two-site anisotropies were assessed experimentally only in few cases [36, 27]. Hence, isotropic NN and NNN interactions are often assumed in the analysis [37], which is further complicated by possible admixture of excited CF levels by magnetic field. Theoretically, ab initio calculations of exchange interactions in Yb delafossites on the basis of DFT electronic structure have so far predicted too small NNN couplings compared to experimental expectations [38, 37].
In this work, we tackle the problem of evaluating realistic spin Hamiltonians for triangular rare-earth systems using DFT+dynamical mean-field theory (DMFT) [39, 40, 41] treating strongly localized rare-earth 4ff states in a quasiatomic (Hubbard-I) approximation [42] to obtain the high-temperature paramagnetic electronic structure, as well as, the 4ff-shell CF splitting and ground-state Kramers doublet. Intersite exchange interactions (IEI) are evaluated from the same paramagnetic electronic structure using the force-theorem in Hubbard-I (FT-HI) method [43]. In contrast to the previous ab initio calculations deriving IEI in triangular rare-earth systems from DFT electronic structure with metallic 4ff bands [38], within our framework the IEI are obtained from realistic self-consistent DFT+HI electronic structure of the target compounds with localized 4ff states. A significant impact of DFT+DMFT charge-density self-consistency on the electronic structure and CF splitting is well documented [44, 45, 46]. We apply this framework to derive realistic spin Hamiltonians for a representative set of Ce- and Yb- triangular compounds, namely: KCeO2, KCeS2, RbCeO2, and YbZn2GaO5 (YZGO). These materials are either experimentally inferred to exhibit conventional magnetic orders (KCeO2 [23] and KCeS2 [22, 24]), a putative DSL (YZGO [33, 47]) or remain to be investigated in detail (RbCeO2, for which no order was found in a preliminary study [25], but a full characterization has not been carried out). We then study the resulting Hamiltonians through exact diagonalization (ED) calculations [48, 49], obtaining numerically exact ground- and low-energy eigenstates resolved by the lattice space group on clusters with sizes up to $36$ spins. This, then, allows us to analyze in detail the nature of the magnetic ground-state (GS) spin-spin correlations for each of the considered compounds. Previous works have already investigated certain regions of the parameter space for these type of Hamiltonians, through classical Monte Carlo studies [50, 51], ED [52] and notably using DMRG [32, 53, 54, 55].
Ab initio method. We employ a charge self-consistent implementation [56, 57, 58] of DFT+DMFT based on the Wien2k linearized augmented-plane-wave (LAPW) full-potential code [59] and ”TRIQS” library [60]. All calculations were carried out using the reported experimental lattice structures [23, 24, 25, 33]. The IEI are calculated by FT-HI approach [43] using the MagInt code [61] analogously to previous applications of this approach to correlated magnetic insulators, see Ref. [62] for a review. The FT-HI method includes all kinetic exchange contributions due to virtual hopping of 4ff electrons. See the Supplemental Material (SM) [63] for more details on our ab initio calculations and the effect of charge-density self-consistency in DFT+HI.
Figure 1: Calculated CF level splitting. The thick red lines are calculated CF levels; the dashed blue lines are CF excitations measured in inelastic neutron scattering (INS) from Refs. [23, 22, 25, 33]. Note that three excitations are seen in INS spectra for the Ce compounds.
KCeO2 KCeS2 RbCeO2 YZGO
$g_{ab}$ This work 1.43 2.00 1.54 3.24
Refs. 2.00a 1.67b, 2.47c 1.46d 3.04e
$g_{c}$ This work 0.17 -0.10 0.30 3.04
Refs. 0.29a 0.58b,0.65c 0.01d 3.44e
Table 1: Calculated values for the in-plane ($g_{ab}$) and out-of-plane ($g_{c}$) gyromagnetic tensor components. Other values are obtained by electron spin resonance (ESR), Refs. [23] (a) ESR and quantum chemistry calculations (QCC) [22] (b); QCC, Ref. [24] (c); INS, Ref. [25] (d); magnetization measurements, Ref. [33] (e).
Crystal-field splitting and ground state. DFT+HI calculations converge to the expected insulating state with $f^{1}$ and $f^{13}$ GS occupancy of the Ce and Yb 4ff shells, respectively (see the SM for the calculated DFT+HI spectral functions). The calculated CF splitting are shown in Fig. 1 and the corresponding Kramers doublets are listed in the SM [63]. We find a particularly large total CF splitting – about 60% of the SO gap (the latter is 0.3 eV in all Ce systems) – for the Ce oxides. Correspondingly, there is a non-negligible admixture $\approx 0.17|7/2;\pm 1/2\rangle$ of the SO excited multiplet to their GS Kramers doublets. The CF splitting in KCeS2 is almost 3 times smaller leading to the corresponding reduction of the $J=7/2$ admixture. No noticeable multiplet mixing is found for YZGO due to its much larger SO gap of 1.37 eV. The dominating contribution of $|M=\pm 1/2\rangle$ to the GS of all Ce systems leads to an in-plane magnetic anisotropy. In contrast, the $|M=\pm 7/2\rangle$ states provide the largest contribution to the GS doublet of YZGO. The corresponding $g$-tensor extracted from the calculated magnetic moment operators $M_{\alpha}=g_{\alpha}S^{\alpha}$, $\alpha=x,y,z$ within the GS Kramers doublet is listed in Table 1. Here we introduced the pseudo-spin $S=1/2$ to label the states within the GS doublet and the corresponding spin operators $S^{\alpha}$. One sees that the in-plane components $g_{ab}=g_{x}(=g_{y})$ largely dominate in all Ce systems. In KCeS2, the out-of-plane component even becomes negative as a result of a near-perfect cancellation between the J=J=5/2 states, with the moment induced by a J=J=7/2 admixture. The $g$-tensor in YZGO is significantly larger and quasi-isotropic.
The experimental determination of the CF splitting and GS in the Ce compounds under consideration is complicated by the appearance of an additional, third, mode in their INS spectra [21, 25, 22], though only two CF excitations are expected in the case of a $J=5/2$ GS. This phenomenon, observed also in other Ce systems, remains so far without commonly accepted explanation, see Refs. [24, 64, 65]. Taking this complication into account, our predictions for the CF levels energies in all compounds under consideration are in good agreement with experiment (Fig. 1), albeit with some systematic underestimation of the splitting by the present theory. Similarly, a good agreement is found for the $g$-tensor (Table 1) in comparison with previous experimental estimates and quantum chemistry calculations. We note that no signal was observed in the electronic spin resonance for the out-of-plane direction in KCeS2 suggesting $g_{c}\ll g_{ab}$ [22].
Effective magnetic Hamiltonian. The IEI within the GS Kramers manifold were calculated by the FT-HI method. Besides IEI, we also included the classical intersite dipole-dipole interaction, which can be important given rather small magnitudes of anisotropic IEI in the target materials, evaluated using the ab initio gyromagnetic tensors (Table 1).
The symmetry of the triangular rare-earth layer constrains the most general coupling within the $S=1/2$ manifold for the NN $ij$ bond along $\mathbf{r}_{0}^{\mathrm{NN}}=[100]||x$ to [29, 54, 66]:
$H_{ij}$ $=\mathbf{S}_{i}^{T}\hat{J}\mathbf{S}_{j}=J(\Delta S_{i}^{z}S_{j}^{z}+S_{i}^{x}S_{j}^{x}+S_{i}^{y}S_{j}^{y})+$ (1)
$2J_{\pm\pm}(S_{i}^{x}S_{j}^{x}-S_{i}^{y}S_{j}^{y})+J_{z\pm}(S_{i}^{z}S_{j}^{y}+S_{i}^{y}S_{j}^{z}),$
where $\mathbf{S}^{T}_{i}=[S^{x}_{i},S^{y}_{i},S^{z}_{i}]$, $\Delta$ is the diagonal XXZ anisotropy, $J_{\pm\pm}$ is the diagonal XY anisotropy and $J_{z\pm}$ are off-diagonal terms. The coupling for the NNN bond along $\mathbf{r}_{0}^{\mathrm{NNN}}=[010]||y$ takes the same form of Eq. (1); we will use the prime to distinguish the NNN terms ($^{\prime},J±±J_{\pm\pm}^{\prime}$ etc.). The coupling for other NN and NNN bonds follows from the symmetry [67, 66], see SM [63] for details.
KCeO2 KCeS2 RbCeO2 YZGO
$J$ (meV) 0.51 0.021 0.36 0.25
$\Delta$ 1.03 1.90 1.05 1.03
$J_{\pm\pm}/J$ 0 0.18 0 -0.055
$J_{z\pm}/J$ 0.10 -1.45 -0.121 0.039
$J^{\prime}/J$ -0.055 -0.023 -0.063 0.003
$\Delta^{\prime}$ 0.70 -0.64 0.67 5.96
$J_{\pm\pm}^{\prime}/J^{\prime}$ 0.009 -0.62 -0.002 2.62
$J_{z\pm}^{\prime}/J^{\prime}$ 0.043 1.00 -0.040 0.073
$J^{{}^{\mathrm{int}}}/J$ 0.011 0.016 0.012
$\theta$ (K) -8.5 -0.43ab -6.2 -3.9ab,-1.9c
$\theta_{\mathrm{Exp}}$ (K) -7.7 -2.8ab -5.1 -5.2ab,-3.8c
Table 2: Calculated intersite coupling including the IEI and dipole-dipole contributions. Experimental Curie-Weiss temperatures $\theta_{\mathrm{Exp}}$ extracted from out-of-plane (c), in-plane (ab), or polycrystalline data are from Refs. [23, 22, 25, 33].
The resulting intersite couplings are listed in Table 2. All systems exhibit AFM NN $J$, as expected; we find $\Delta\gtrsim 1$ everywhere. The NNN interactions are ferroic in all Ce systems and antiferroic in YZGO. There is an essential difference between the Ce and Yb systems in the nature of NNN interactions: in the former, the dipole-dipole interactions are negligible compared to the exchange. In YZGO, due to small NNN IEI and large $g$ factors, the anisotropic dipole-dipole contribution dominates over an almost isotropic NNN exchange. As a result, the total NNN couplings in YZGO are very anisotropic ($\Delta\approx 6$) with $J_{xx}^{\prime}(=J^{\prime}+2J^{\prime}_{\pm\pm})=0.020J$, $J_{yy}^{\prime}(=J^{\prime}-2J^{\prime}_{\pm\pm})=-0.013J$, and $J_{yy}^{\prime}=\Delta J^{\prime}=0.019J$. Hence, though the NNN interactions in YZGO are small, they are not as negligible as may seem from the $J^{\prime}/J$ ratio listed in Table 2. The effect of a strong NNN anisotropy on the phase stability on the triangular lattice has not been studied so far. Finally, the coupling $J^{\mathrm{int}}$ for the shortest interlayer bond is about 1%-1.5% of $J$ in all oxide systems (we could not extract it in KCeS2 due to the general smallness of IEI in this system).
Figure 2: Spin structure factors and low-energy spectrum for the YZGO ($a,b$) and KCeS2 ($c,d$) compounds in a $36$-site cluster. The isotropic structure factor $\mathcal{S}\left(\bm{k}\right)$ in $a)$ is peaked at the $K$ points, indicating $120^{\circ}$ order, consistent with low-energy excitations at $K_{0}.\text{A}$ and $K_{1}.\text{B}$. The peak at $M_{2}$ of $\mathcal{S}^{yy}(\bm{k})$ in $c)$ indicates stripe-$\perp$ order, consistent with low-energy excitations at the $M$ points in $d)$.
To assess the overall agreement in magnitude with experiments, we calculated high-temperature susceptibility by solving (1) in mean-field in an applied field using the McPhase package [68] and an in-house module. For all oxide systems we find quite good agreement of the calculated Curie-Weiss temperatures $\theta$, extracted from the averaged susceptibility $(2\chi_{ab}+\chi_{c})/3$, with available experimental data (Table 2). In the case of KCeS2, theoretical intersite couplings are severely underestimated. This is also apparent from comparison of our result with the estimates of Ref. [69] that were obtained by quantum chemistry calculations and subsequently adjusted to agree with experimental INS. 4ff kinetic exchange captured by FT-HI does not, apparently, give the leading contribution to the IEI in this sulfide.
Low-energy spectra and ground-state correlations. We now investigate the GS phase diagram of the effective magnetic Hamiltonians introduced in Eq. (1) with ED using the XDiag library [49, 70]. We obtain symmetry-resolved excitation spectra employing translational symmetries and the lattice inversion symmetry without spin rotation on cluster sizes $N=12,32,36$, see the SM [63] for the employed lattice geometries. The irreducible representations (irreps) are represented by $\bm{k}.\rho$, where $\bm{k}$ labels the momentum and $\rho$ labels the inversion irrep, i.e. either $\rho=\text{A}$ for the even irrep or $\rho=\text{B}$ for odd irrep. We note that the additional point group symmetries are not employed, as an implementation of combined lattice-spin rotations is not readily accessible. We also obtained the GS static spin structure factors,
$\mathcal{S}^{\alpha\alpha}\left(\bm{k}\right)=\dfrac{1}{N}\sum_{n,m}e^{i\bm{k}\cdot\left(\bm{r}_{n}-\bm{r}_{m}\right)}\left\langle S^{\alpha}_{n}\cdot S^{\alpha}_{m}\right\rangle,$ (2)
where $\alpha=x,y,z$, $N$ denotes the number of lattice sites and $\bm{r}_{n}$ the position of the $n^{\rm th}$ spin. Furthermore, we denote by $\mathcal{S}\left(\bm{k}\right)=\sum_{\alpha}\mathcal{S}^{\alpha\alpha}\left(\bm{k}\right)$, the isotropic spin structure factor.
In panels $a)$ and $b)$ of Fig. 2, we show $\mathcal{S}\left(\bm{k}\right)$ and the low-lying energy spectrum for the model parameters of YZGO (see table 2) on the $N=36$ cluster. Both the organization of the spectrum and the spin structure factor are consistent with the system realizing the $120^{\circ}$ AFM phase. The GS transforms under the $\Gamma.\text{A}$ irrep, while the first excited states (nearly degenerate) belong to the $\Gamma.\text{B}$ and $K.\text{A}$ irreps. These symmetry quantum numbers match the expected signatures of the $120^{\circ}$ AFM phase [71, 72], see the SM [63] for further details on how these quantum numbers are predicted.
A key signature of a DSL is a gapless singlet mode at the $X_{0}$ point in the first Brillouin zone (FBZ), corresponding to the singlet monopole excitations [19]. Although in the present case we cannot identify these excitations as singlets due to the absence of spin SU(2) symmetry, they still occur at the same momentum. In contrast, the excitation energy at $X_{0}$ is significantly larger compared to the putative DSL of the $J_{1}$-$J_{2}$ model [11], rendering this scenario unlikely. Furthermore, $\mathcal{S}(\bm{k})$ exhibits a pronounced peak at crystal momentum $K$, in agreement with the known ordering wave vector of the $120^{\circ}$ AFM state [2, 73]. Analogously, we also observe spin correlations and a low-energy spectrum consistent with a $120^{\circ}$ state for the effective Hamiltonians of the KCeO2 and RbCeO2 compounds, see the SM [63].
Figure 3: $a)$ and $d)$ - Approximate phase diagram of the effective magnetic Hamiltonian, inferred from the quantum number organization of the low-energy spectrum. Dark red marks the $120^{\circ}$ AFM phase, dark blue the stripy-$\parallel$ order, pale purple the stripy-$\perp$ order, and orange the regime with first non-zero momentum excitation in the $X_{1}.\text{A}$ irrep, suggesting a possible DSL phase. A typical classical ordering for each stripe phase is shown in the sketch in $d)$. The remaining panels show the static spin structure factor at selected momenta in the FBZ (see inset in $a$)): panel $b)$ and $e)$ at $Z$, while $c)$ and $f)$ show the average over the three inequivalent $M$ points. In $b)$ we identify the parameters compatible with the YZGO system. Parameters: $\Delta=1.03J$, $J^{\prime}=0.02J$ in a)a)$c)$, and $J_{z\pm}=0$ in d)d)$f)$.
On the other hand, the spectrum and correlations for the KCeS2 system indicate a collinear stripy antiferromagnetic GS, as shown in panels $c)$ and $d)$ of Fig. 2. The low energy spectra reveals an approximate GS degeneracy between the $\Gamma$.A, and $M.\text{A}$ irreps, consistent with the stripy AFM state [72]. Accordingly, we observe a pronounced peak in $\mathcal{S}^{yy}(\bm{k})$ and $\mathcal{S}^{zz}(\bm{k})$ at momentum $M_{2}$The peak appears at only one $M$ point because slight numerical rounding in the ab-initio couplings weakly breaks the Hamiltonian’s $C_{3}$ symmetry.. As noted [50], the symmetries of Eq. (1) allow for two distinct stripe orders, differing by how they transform under the combined lattice-spin rotation around the triangular lattice axis directions. The first aligns the spins along the direction of the stripe, we call it stripe-$\parallel$ (stripe-parallel) (this is also named stripe-$x$). It is invariant under the $C_{2}$ axis rotations, as the spins align with it. In the second, the spins align perpendicular to the stripe direction, and thus transforms non-trivially under the $C_{2}$ axis rotation. We name it stripe-$\perp$ (stripe-perpendicular) (also known as stripe-$yz$). Thus, the (in)variance under this symmetry constitutes the distinguishing feature between the two states . A representative configuration of each stripe type is shown in Fig. 3 $d)$.
The stabilized stripe is oriented along the $x$-direction of the lattice, so, the peak in the $\mathcal{S}^{yy}(\bm{k})$ at $M_{2}$ indicates the realization of the stripe-$\perp$ state, see the SM for further details. Remarkably, this is in complete agreement with experiments on KCeS2 [24].
[…]
We perform extensive ED calculations on the $32$-site cluster and use the $12$-site cluster to corroborate the phase diagram obtained, whose results and discussion are left to the SM [63].
We obtain an approximate phase diagram (Fig. 3 $a)$ and $d)$) based on the quantum numbers of the spectrum (check the SM). We begin by fixing $J^{\prime}=0.02J$ (a value compatible with the largest matrix elements of the NNN coupling $\hat{J^{\prime}}$ in YZGO; see Table 2) and study the phase diagram as a function $J_{\pm\pm}$ and $J_{z\pm}$. This corresponds to the panels ($a,b,c$) in the upper row of Fig. 3. In this case, the phase diagram consists of three distinct magnetically ordered states, a $120^{\circ}$ AFM region, a stripe-$\parallel$ phase, and a stripe-$\perp$ phase, which correspond to the dark red, dark blue, and pale purple regions shown in panel $a)$, respectively.
As in the 32-site cluster, the $K$ momentum is not resolved, $\mathcal{S}(\bm{k})$ peaks at nearby wave vectors $Z_{0}$ and $Z_{1}$ in the $120^{\circ}$ AFM phase, as shown in Fig. 3 $b)$. This reflects the frustrated realization of $120^{\circ}$ AFM order in this cluster.
Upon tuning both $J_{\pm\pm}$ and $J_{z\pm}$, $\mathcal{S}(\bm{k})$ develops a peak at the $M$ points, signaling the onset of the stripe phase (see panel $c)$) We find that along the $J_{z\pm}=0$ line, negative values of $J_{\pm\pm}/J$ favor a stripe-$\parallel$, whereas positive values stabilize a stripe-$\perp$ state. The coupling $J_{z\pm}$ further extends the stripe-$\perp$ region into negative $J_{\pm\pm}/J$ values, which agrees with previous studies. In certain parameter regimes, the peak in $\mathcal{S}(\bm{k})$ appears at both $M_{0}$ and $M_{1}$, while in others it shifts to $M_{2}$. This shift does not signal a transition between stripe phases, but originates from the absence of $C_{3}$ symmetry in the $32$-site cluster, which renders the three $M$ points inequivalent (unlike in the $12$-site cluster). Consequently, the lattice geometry lifts the degeneracy among the six stripe configurations. To account for this, in Fig. 3 $c)$ we plot $\mathcal{S}(\bm{k})$ averaged over the three $M$ points,
$\tilde{\mathcal{S}}(M)=\frac{\mathcal{S}(M_{0})+\mathcal{S}(M_{1})+\mathcal{S}(M_{2})}{3}.$ (3)
Distinguishing the two stripe types requires examining additional spin-spin correlators, which we do in the SM. We observe in panel $b)$ that the couplings determined for the YZGO system place it within the $120^{\circ}$ AFM region as expected.
We set $J_{z\pm}=0$ and study the phase diagram as a function of $J^{\prime}$ and $J_{\pm\pm}$, identifying four distinct regions, as shown in Fig. 3 $d)$. For $J^{\prime}\in[0,0.05]J$, the GS is in a $120^{\circ}$ AFM state, confirmed by the peak of $\mathcal{S}(\bm{k})$ at the $Z$ points (panel $e)$). This state persists up to $J_{\pm\pm}\sim 0.1J$. Increasing $J_{\pm\pm}$ induces stripe order, whose orientation depends on the sign of $J_{\pm\pm}$: negative values yield stripe-$\parallel$, while positive values stripe-$\perp$ GS. This is corroborated by the peak of $\mathcal{S}(\bm{k})$ at the $M$ points as shown in panel $f)$.
For $J^{\prime}\sim[0.05,0.16]J$ and $J_{\pm\pm}\sim[-0.1,0.1]J$, the first non-zero momentum excitation belongs to the $X_{1}.\text{A}$ irrep (see SM), corresponding to the orange region in panel $d)$. As discussed, the excitation at momentum $X$ is particularly noteworthy, as it is interpreted as a signature of the singlet-monopole of a DSL [19, 20], predicted to appear in the low-energy spectrum of the $J_{1}$$J_{2}$ triangular-lattice antiferromagnet in the spin-liquid regime [11]. An alternative explanation for such a low-lying excitation is a 12-site valence-bond solid, for which consistent correlations have also been observed in the spin-liquid regime [11]. We find that this region is stabilized only for small values of $J_{\pm\pm}\sim\pm 0.1J$, while larger values drive the system towards the stripe phase. This analysis indicates that the previously reported quantum paramagnetic phases in earlier DMRG studies [54] (identified even in the absence of NNN coupling) might not correspond to a DSL phase. In the SM, we provide additional results of the energy spectra, which reveal that the $X.\text{A}$ states persist at higher energies, whereas the low-lying excitations belong to the tower of states characteristic of the $120^{\circ}$ AFM order. While these results do not definitively exclude other paramagnetic states, the sharp $K$-point peak in the spin structure factor (computed also for an $N=36$ cluster at $J^{\prime}=0$, $J_{z\pm}=0.3J$, and $J_{\pm\pm}=-0.08J$) strongly supports $120^{\circ}$ AFM order. Although a DSL may exhibit similar quasi–long-range signatures, which finite-size ED cannot unambiguously distinguish from true long-range order, our spectral analysis offers little evidence for such a phase at this parameter point.
Conclusions – We have derived from first principles the crystal-field splitting and and intersite exchange interactions within the Kramers GS manifold for a set of rare-earth triangular materials, including two systems (KCeO2 and KCeS2) that order magnetically, as well as two putative QSLs (RbCeO2 and YZGO). The calculated CF splitting and CF GS $g$-factors are in good agreement with experiment. The mean-field Curie-Weiss temperature evaluated from the resulting low-energy Hamiltonian agrees well with experiment for all three oxide systems but is severely underestimated for KCeS2.
Solving the resulting low-energy Hamiltonian with ED, we find evidence for the $120^{\circ}$ order in all systems apart from KCeS2, where a stripe-$\perp$ state is instead stabilized, consistent with experimental observations. We find strong evidence for a magnetic $120^{\circ}$ order in the two DSL candidates. From the extended phase diagram of their Hamiltonian, we concluded that a stronger antiferromagnetic NNN coupling is required to realize the DSL phases in comparison to those present in the compounds.
We notice that a very recent neutron-scattering study [75] reports significant site disorder in YZGO as evidenced by broadening of crystal-field excitations and small random crystalline distortions. The effect of site disorder was discussed in the context of a well-studied similar system, YbMgGaO4, where the resulting randomization of intersite exchange was predicted to mimic a QSL state [32]. Our results together with the observations of Ref. [75] suggest that the same phenomenon may be at play in YZGO.
Our calculated intersite exchange interactions include all $f$-$f$ kinetic exchange contributions (in particular, superexchange). However, in rare-earth semiconductors there can be another type of $f$-$f$ intersite exchange generated through 4ff-5dd hopping and the inter-shell Hund’s rule interaction [76]. This mechanism could become important in the case of very weak superexchange like in KCeS2, where our calculations apparently reproduce the structure of intersite exchange but not its magnitude. Its impact on the NNN exchange $J^{\prime}$ also needs to be evaluated, especially in the light of our present predictions for $J^{\prime}$ to be too small to stablize a DSL phase in YZGO. Further studies to evaluate the importance of such exchange mechanisms in triangular rare-earth compounds are worthwhile.
Acknowledgments – The exact diagonalization calculations have been performed using the XDiag library [49]. L.V.P. is grateful to Sidhartha Dash, Dario Fiore Mosca, and Antoine Georges for useful discussions. L.V.P acknowledges the support by CNRS through the Tremplin@Physique 2025 program and by the CPHT computer team. A.W. acknowledges support by the German Research Foundation (DFG) through the Emmy Noether program (Grant No. 509755282).